define a variable and write an equation for the situation. then solve.
suppose you are driving to visit a friend in another state. you are driving 50 miles per hour. you must drive 600 miles total. if you have already driven 375 miles, how long will it take you to reach your destination?

Answer by nerdybill(5409)

suppose you are driving to visit a friend in another state. you are driving 50 miles per hour. you must drive 600 miles total. if you have already driven 375 miles, how long will it take you to reach your destination?
.
Applying "distance formula":
d = rt
where
d is distance
r is rate or speed
t is time
.
Solving for 't':
t = d/r
.
t = (600-375)/50
t = 225/50
t = 4.5 hours